three resistors are wired in series with a 22v battery. The resistances are 22.5 omega 33.6 omega and 9.9 omega. what is the voltage across the 9.9 omega resistor
Answer #1

Voltage (v) = 22v

Total resistance in series is Req = 22.5 + 33.6 + 9.9 = 66 omega

So current in the circuit is ( I )= voltage / total resistance = 22/66 = 0.333A

So voltage across the resistor 9.9 omega is (v) = current * resistance = 0.333 * 9.9

Voltage = 3.2967 volt = 3.3 volt

plz upvote